Solve the system of linear equations. Write your answer as an ordered pair (x, y, -3).
x - y + 2z = 22
3y - 8z = -9
z = -3
(x, y, -3) = ?

Answer by Alan3354(69443)

Solve the system of linear equations. Write your answer as an ordered pair (x, y, -3).
x - y + 2z = 22
3y - 8z = -9
z = -3
-----------
Sub for z in the 1st 2 eqns:
x - y - 6 = 22
3y +24 = -9
---------
Solve the 2nd eqn for y.
Then solve for x.
-----
Write it any way you like.